OpenStudy (anonymous):

When I try solving tan 5pi/6 I keep getting 1/-sqrt3

OpenStudy (campbell_st):

that's correct... as this is a 2nd quadrant angle where tan is negative to check find \[\tan (\frac{\pi}{6})\] from the exact value triangle  and then \[\tan(\frac{5\pi}{6}) = - \tan(\frac{\pi}{6})\] hope it helps

OpenStudy (anonymous):

but it isn't one of my answer choices. the right answer is -sqrt3/3 and i don't understand how that is @campbell_st